A mass of 2 kg on a horizontal surface with coefficient of friction equal to 1/2 is connected to a second mass of 4 kg over a frictionless pulley. If g = 10 ms-2, calculate the acceleration of the system when released
1
Expert's answer
2020-06-25T09:35:51-0400
"(m+M)a=Mg-\mu mg=(M-\mu m)g"

"(2+4)a=(4-(0.5)2)9.8"

"a=4.9\frac{m}{s^2}"